|up||API Server||100.000 %|
|up||Client Area||99.869 %|
|up||Dallas Network 02||100.000 %|
|up||Dallas Nwtwok 01||100.000 %|
|up||DAX/Main Server||100.000 %|
|up||Lenoir Network 01||100.000 %|
|up||NYC Network 01||99.888 %|
|up||Vint Hill Network 01||100.000 %|
Date: 09/15/2020, @ 14:30 (02:30 PM) EST Services Affected: Backup01 and VPN VPN-NYC 01 Current Status: Resolved Resolved - The issue has been resolved. In Progress – After looking at the problem we have found out that we have a bad switch. We are configuring and...
Hello, Maintenance is required for SH101. The host will be rebooted, and a number of critical updates will be installed. These include general performance improvements, enhancements to networking throughput, and a BIOS upgrade intended to address the primary cause of...
Date: 07/11/2020, @ 14:30 (02:30 PM) EST Services Affected: Mail01 and Client Area Current Status: Resolved Resolved - This issue has now been resolved. Monitoring - The data center has reported that this issue has now been resolved. We’ll keep an eye on it and send...
The server has been under attack and due to the size the mitigation has nulled the IP. We're working with the DC to get things back online in a safe manner. Resolved - The issue has now been resolved. If you have any questions, please send a support ticket....
Original Report: At 01:50 AM EDT, we were notified by monitoring software that VPS Node Node01 was not available. We are further investigating the issue now and once we have more information, it will be posted here. Update at 03:15 AM EDT 6/11/2020: We are unable to...